OpenStudy (mini):

fxy if f(x,y) = (xy)^(1/2)-1/(x^2*y)

OpenStudy (yuki):

what would you want to do?

OpenStudy (mini):

its double der

OpenStudy (mini):

x then y

OpenStudy (yuki):

ok, so you are finding \[f_{xy}\]

OpenStudy (mini):

yes

OpenStudy (mini):

its just really long wanted to see someone else do it lol

OpenStudy (anonymous):

derivate everything to respect x and then derivate that answer to respect y

OpenStudy (yuki):

let me re-write the eqn first, it's tough to see it (xy)^(1/2)-1/(x^2*y) \[{(xy)^{1/2} \over x^2y}\]

OpenStudy (yuki):

am I right?

OpenStudy (mini):

already know that brain but thank you

OpenStudy (mini):

ill write what i have first

OpenStudy (mini):

ill write what i have first\[.5(xy)^{-1/2}y+(x ^{2}y)^{-2}2xy\]

OpenStudy (anonymous):

hmm.. it's going to be a quotien rule with two product rules. sorry but i'm kinda lazy to actualy write it

OpenStudy (mini):

quotient? i just did two product rules...

OpenStudy (yuki):

what you should notice first is that the equation is equivalent to \[\sqrt{1 \over yx^3}\]

OpenStudy (yuki):

this is a lot easier to find the derivatives

OpenStudy (mini):

i do chain rules and product, u dont have to use quotient

OpenStudy (yuki):

Mini, just to let you know the quotient rule makes these problems far more faster, although I know that it is a pain. so I will just do this one with quotient rule and you try to confirm it with product rule, ok?

OpenStudy (mini):

well product and quotient are the same thing though lol, its like comparing addition of negatives and subtraction

OpenStudy (anonymous):

you are missing a -1 on the top part -> (xy)^(1/2)-1/(x^2*y): \[(\sqrt{xy}-1)\div(x^{2}y)\]

OpenStudy (yuki):

to find f_x, I will write f as follows \[{1 \over \sqrt y} * {1 \over x^{3/2}}\]

OpenStudy (mini):

dont see what im missing i might of had a typo but my dx is + to the neg 2

OpenStudy (yuki):

Brain, nice catch, I was going to solve the problem not knowing that lol

OpenStudy (anonymous):

haha yea, sorry i was trying to type on the "Equation" thing it took me a while lol

OpenStudy (yuki):

then f = \[\sqrt{xy} - {1 \over x^2y}\]

OpenStudy (mini):

yuki i honestly have no idea what you are doing, im changing everything to exponents then doing chain and products, it might be "slower" but its faster for myself

OpenStudy (mini):

i mean chain is pretty simple, bring to the front, copy, minus 2, der of inside

OpenStudy (yuki):

then f_x is \[{\sqrt{y} \over 2\sqrt{x}}+{2 \over yx^3 }\]

OpenStudy (mini):

minus 1 even

OpenStudy (mini):

yeah thats not what i got .5(xy)^(-1/2)y

OpenStudy (mini):

well the first time with respect to x

OpenStudy (yuki):

now f_xy is \[{1 \over 4\sqrt{xy}} - {2 \over x^3y^2}\]

OpenStudy (anonymous):

ok. Mini, what i got so far was: Fx = \[(-2(\sqrt{xy}-1))\div(x^{3}y)\] Fxy = \[(2\times(\sqrt{xy}-1))\div(y^{2}x ^{3})\]

OpenStudy (mini):

\[.5(xy)^{-1/2}y+(x ^{2}y)^{-2}2xy\]

OpenStudy (mini):

then product x2

OpenStudy (mini):

for respect to y

OpenStudy (yuki):

I am 100% sure with my answer

OpenStudy (yuki):

I heard product rule, but you don't have to use it anywhere. at least I didn't

OpenStudy (mini):

for respect to y\[-.25(xy)^{-3/2}xy+.5(xy)^{-1/2}-2(x ^{2}y)^{-3}2x ^{3}y+(x ^{2}y)^{-2}2x\]

OpenStudy (yuki):

Mini, when you find the partial derivative of x, y is not considered a function of x, so you do not use the product rule

OpenStudy (mini):

so its not fx*fy

OpenStudy (yuki):

so for the same reason we don't do chain rule as well

OpenStudy (anonymous):

no, you dont multiply both derivatives

OpenStudy (mini):

but why is it not in this case? its with respect to y

OpenStudy (anonymous):

you take the derivate to respect x first to get Fx, then you take the derivate to respect y for the new functions whch will turn into Fxy

OpenStudy (yuki):

f_xy means \[f_{xy} = d/dy(d/dx (f))\]

OpenStudy (anonymous):

yes

OpenStudy (mini):

yes i know that much but if we are talking in terms of y, y isnt .25(xy)^(-3/2)y considered two?

OpenStudy (mini):

like y times the other part

OpenStudy (mini):

thats just like saying in terms of x(3-x)

OpenStudy (yuki):

so you take the derivative of f in terms of x while y is constant and derivative of f_x in terms of y while x is constant

OpenStudy (mini):

its the same thing if instead you do x^3, as x times x^2, is it not?

OpenStudy (mini):

the last part is in terms of y, which applies to both parts so im still missing why i cant do the product once i did fx

OpenStudy (mini):

rather respect to y

OpenStudy (mini):

its the same thing if instead you do x^3, as x times x^2, is it not?

OpenStudy (mini):

that question still holds

OpenStudy (anonymous):

Mini, Are yu asking why can't you multiply fx and fy separatly?

OpenStudy (yuki):

Mini, was my f(x,y) correct? I'm wondering because you don't seem to have any interest in what I said.

OpenStudy (anonymous):

like: Fx * Fy

OpenStudy (mini):

im asking why i cant just do \[.5(xy)^{-1/2}y+(x ^{2}y)^{-2}2xy\]

OpenStudy (mini):

and to the product on that, and i dont know yuki, im not worried about the answer im worried about the concept

OpenStudy (mini):

not trying to be a wingspan, plus i dont even have the answer haha

OpenStudy (yuki):

Mini, because y is not a function of x. You do not do product rule with partial derivatives because in partial derivatives, other variables are constant

OpenStudy (yuki):

so let me give you an example

OpenStudy (mini):

but why is that not the same as x*x^2 = x^3

OpenStudy (yuki):

let's say f(x,y) = \[3x^2y\]

OpenStudy (mini):

and separating those and doing one by product and one by chain

OpenStudy (mini):

6xy

OpenStudy (mini):

then 6x

OpenStudy (yuki):

f_x would be yes, 3y*(x^2)' = 6xy

OpenStudy (yuki):

f_y would be 3x^2

OpenStudy (yuki):

f_xy and f_yx must be the same, so let's check f_xy = 6x(y)' = 6x f_yx = 3(x^2)' = 6x yay

OpenStudy (mini):

well yeah i follow that, thats easy lol

OpenStudy (anonymous):

There's no need to do chain rule in this problem

OpenStudy (yuki):

If you would like me to answer your question why (x^3)' and (x*x^2)' are not the same they actually are

OpenStudy (mini):

ah there might be no NEED, but can i do it that way

OpenStudy (mini):

its my train of thought, thats all i can say

OpenStudy (mini):

and i know they are the same, thats what im basing my principle on just doing the product rule on

OpenStudy (yuki):

(x*x^2)' is x'*x^2 + (x^2)'*x =x^2+2x*x =3x^2

OpenStudy (mini):

\[.5(xy)^{-1/2}y+(x ^{2}y)^{-2}2xy\]

OpenStudy (mini):

thats my in terms of x

OpenStudy (yuki):

okay now I think we are on the same page, so you are trying to use the product rule for this problem anyways, am I right ?

OpenStudy (mini):

my question still stands why i cant do the product rule on this

OpenStudy (mini):

yes!!!

OpenStudy (anonymous):

MIni, you can use the product rule

OpenStudy (mini):

and that equation is now going to be derived in terms of y

OpenStudy (yuki):

okay, gotcha. just to let you know that is very unorthodox, and I haven't tried it yet, so give me some time to work it out first.

OpenStudy (mini):

haha, thats what my professor said, i know he loves grading my papers but i see everything as addition and multiplication

OpenStudy (anonymous):

hehe okay, yea, you can do product rule but yu still have to do another quotien rule

OpenStudy (anonymous):

you will have to do two product rules inside a quotien rule

OpenStudy (yuki):

mini, I want to know how you will consider f(x,y)

OpenStudy (mini):

where? once i do the product for \[.5(xy)^{-1/2}y+(x ^{2}y)^{-2}2xy\]

OpenStudy (mini):

on both sides isnt it just two products?

OpenStudy (anonymous):

isn't the main equation a fraction?

OpenStudy (mini):

no, i changed them into multiplication

OpenStudy (anonymous):

can you type the main equation and your "changed" equation? please

OpenStudy (mini):

sqrt(xy)-1/(x^2*y)

OpenStudy (anonymous):

yes that's the main, now what did yu change that for?